Systems Staff Meeting Notes 1/8/01

Systems Staff Meeting Notes
January 8, 2001

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bill advises that we check the list of pager numbers in the Systems Folder. His has changed recently. The new number is noted there.
  2. Need to add to the Weblist that NT8 & NT2 are not to be used for storage.
  3. Mil Willis is now working in Special Collections.
  4. Construction in the back loading dock area is still ongoing. No schedule for completion is known.
  5. Tape disposition - Martha is supposed to meet with Dolores Jenkins and Jan Swanbeck to discuss tape removal. Dolores still wants to speak with Ray Jones about it. Debra will have NERDC check on the use status of tapes we have over there. Status on pending server room: networking- received quote from Digital, no start date yet; electrical - John Humphrey is aware of our requirements.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 54 Dells to be delivered tomorrow (108 boxes). 24 should be delivered to MSL, 30 to Systems. A distribution list for the 30 has been set. The Dell monitors have been coming in with the speakers installed.
  2. Erich's Polaroid slide scanner is still sitting up in DLC; it has not yet been returned. It was not ordered through us - it is not what he requested.
  3. 40 barcode reader wands received from ScanTech.
  4. No ship date received yet for the big server.
  5. Rack quote received, but not ordered yet.
  6. Two-headed Exchange server ordered. Power supplies ordered with the servers.
  7. Room 2 is getting very full. Need to call Rob Roberts.
  8. Ann Lindell requested a laptop for checkout that would be able to run a Net Agent demo when she goes to ALA. Geoffrey is preparing a Gateway Solo for her. It has 240 RAM, will give her a Xircom card. Installing IE 5.0 not 5.5 will work best for Net Agent.
  9. 3 Olympus Digital cameras have been received. One for Special Collections - Jeffrey Barr says it is not sufficient for them - he had specifically requested a different brand with different specs. Barbara Oliver currently has his Olympus - we need to retrieve it from her so that we can possible use it for staff checkout.
Liaisons
  1. Need to get staff to stop sending multi-megabyte attachments in email. Alternative would be to set up a Web page. Some people may need basic guidelines - the liaison in their area should be able to handle it.
Networking
  1. DLC - was experiencing intermittent trouble accessing NT12. The trouble cleared up mid-day on Friday. The cause of the problem is still unknown. The error message they got was that it was unable to locate the resource name. Could be a WINS problem or update on LM Host file. Need to check the error statistics on the switch.
  2. The disappearance of 2 machines from DLC & 1 from Resource Services has brought up the issue of security. Some suggestions include hard drive locks and tracking software (CyberAngel). Systems can report the theft to Gateway; eventually those machines may be returned to them for service. Need to remind the staff to be more security conscious.
  3. Phones - generally, East seems to be working & all lines and extensions are up to date. Need to call Telecommunications to verify that they have a current backup of the system. Debra will call in a work order on it.
  4. Xerox/NT conversion started last Thursday. Right now it is set up in one section of the public area in West. We'll leave it running for a week then set it up over in Music. We are still running both the Novell & NT servers in West. Need to ensure that IPX has been removed from the test NT machines in West. Marvin has complained that it is not working to his satisfaction on machines running Win'95. Those machines need to be upgraded to Win'98.
  5. Wireless parts are on order. There is a wire hanging from the ceiling of the 1st floor closet that Will thinks probably indicates that Facilities is working on the wiring for the wireless system. Bill will get an update regarding the campus wireless projects at the NTAC meeting this afternoon.
Software Projects Status
  1. Spine Label is working fine.
  2. Net Agent - Winston set up the database. He set up a limit on the size and it hit the top of the transaction log. Backing up the log will clear it out. Winston called Net Agent, but they have no standard on the transaction log size. Adding the backup should take care of the problem.
  3. NT4 - talked about it for clustering. Winston thinks it has lots of space, but will verify. Mission critical: the database and email.
DP Coordinator
  1. Maps' CD ROM conversion - will be an ongoing project for quite awhile. It seems that some of the images to CDs seem to be going on the Web server. Need to make sure that NT8 is not being used for storage. May want to get a hierarchy for permissions for public folder creation in Exchange.
  2. Students have returned - new front desk student is Kristen Dansby. New Olympus digital camera for Systems - Winston checked it out this weekend - was wondering where we'll store it. We'll store it in the vault - should probably move the laptops out of the tape vault and into the locked vaults as well. Debra will need to get the vault combinations. Bill says the camera is for Systems use only, not checkout.


Back to the Systems Homepage

Last updated January 23, 2001
by Debra Harris
debharr@mail.uflib.ufl.edu